a father is 30years older than his son. 3years ago he was 3times as old as his son. How old is the father now?
Answer by Cromlix(4381)

Father = 30 + Son
F = 30 + S
3years ago he was 3times as old as his son.
F - 3 = 3(S - 3)
Replace F with 30 + S
30 + S - 3 = 3(S - 3)
27 + S = 3S - 9
2S = 36
S = 18
Father is 48 years old
Son is 18 years old.
Hope this helps
:-)